PHOTODISSOCIATION SPECTROSCOPY AND DISSOCIATION DYNAMICS OF TiO^{+}(CO_{2})

\underline{MANORI PERERA^{a},} RICARDO B. METZ^{b}, Department of Chemistry, University of Massachusetts Amherst^{b}, Amherst, MA.

TiO^{+}(CO_{2}) is produced by reaction of laser-ablated titianium atoms with CO_{2} and subsequent clustering, supersonically cooled and its electronic spectroscopy characterized by photofragment spectroscopy, monitoring loss of CO_{2}. The photodissociation spectrum consists of a vibrationally-resolved band in the visible, with extensive progressions in the covalent Ti-O stretch (952 cm^{−1} vibrational frequency and 5 cm^{−1} anharmonicity), and in the TiO^{+}-(CO_{2}) stretch (186 cm^{−1}) and rock (45 cm^{−1}). The band origin is at 13918 cm^{−1}, assigned using titanium isotope shifts, and the spectrum extends to 17350 cm^{−1}. The excited state lifetime decreases dramatically with increasing internal energy, from 1100 ns for the lowest energy band (v_{T iO}=0), to <50 ns for v_{T iO}=3. The long photodissociation lifetime substantially reduces the photodissociation quantum yield at low energy, likely due to competition with fluorescence. Electronic structure calculations help to assign the spectrum of TiO^{+}(CO_{2}) and predict allowed electronic transitions of TiO^{+} in the visible, which have not been previously measured. Time-dependent density functional calculations predict that the observed transition is due to B, ^{2}Π ← X, ^{2}Δ in the TiO^{+} chromophore, and that binding to CO_{2} red shifts the TiO^{+} transition by 1508 cm^{−1}, and lowers the Ti-O stretch frequency by 16 cm^{−1}. Combining the computational and experimental results, the ^{2}Π state of TiO^{+} is predicted to lie at T_{0} = 15426 cm^{−1}, with frequency ωe = 968 cm^{−1} and anharmonicity ωexe = 5 cm^{−1}. The calculations also predict that there is only one low-lying ^{2}Σ state of TiO^{+}, contrary to conclusions derived from photoelectron spectroscopy of TiO. Prospects for astronomical observation of TiO^{+} via the ^{2}Π-^{2}Δ transition are also discussed.

a

aDepartment of Chemistry, University of Illinois at Urbana-Champaign